OpenStudy (anonymous):

what is the solution of systems of equations? y+2x=2 y+4x=0

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@nina252000 for question one (y + 2x = 2) y + 2x = 2 2x + y + -y = 2 + -y 2x = -y + 2 2x/2 = -y + 2/2 x = -1/2y + 1

OpenStudy (anonymous):

@nina252000 for question two y + 4x = 0 y + 4x = 0 4x + y + -y = 0 + -y 4x = -y 4x/4 = -y/4 x = -1/4y